# CliBench: Multifaceted Evaluation of Large Language Models in Clinical Decisions on Diagnoses, Procedures, Lab Tests Orders and Prescriptions

### v1

@@ Abstract
The integration of Artificial Intelligence (AI), especially Large Language Models (LLMs), into the clinical diagnosis process offers significant potential to improve the efficiency and accessibility of medical care. While LLMs have shown some promise in the medical domain, their application in clinical diagnosis remains underexplored, especially in real-world clinical practice, where highly sophisticated, patient-specific decisions need to be made. Current evaluations of LLMs in this field are often narrow in scope, focusing on specific diseases or specialties and employing simplified diagnostic tasks. To bridge this gap, we introduce CliBench, a novel benchmark developed from the MIMIC IV dataset, offering a comprehensive and realistic assessment of LLMs’ capabilities in clinical diagnosis. This benchmark not only covers diagnosis from a diverse range of medical cases across various specialties but also incorporates tasks of clinical significance: treatment procedure identification, lab test ordering and medication prescriptions. Supported by structured output ontologies, CliBench enables a precise and multi-granular evaluation, offering an in-depth understanding of LLM’s capability on diverse clinical tasks of desired granularity. We conduct a zero-shot evaluation of leading LLMs to assess their proficiency in clinical decision-making. Our preliminary results shed light on the potential and limitations of current LLMs in clinical settings, providing valuable insights for future advancements in LLM-powered healthcare.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2024; Anil et al., 2023) have opened up new possibilities. They have demonstrated remarkable capabilities in text understanding, reasoning, and generating proper responses in various domains (Ma et al., 2024b), including the medical field (Singhal et al., 2023a, b; Ma et al., 2023; Xu et al., 2023). Specifically, LLMs have been thoroughly evaluated and shown to excel in medical licensing examinations (Nori et al., 2023b; Gilson et al., 2023) and medical knowledge QA benchmarks (Singhal et al., 2023b; Nori et al., 2023a; Toma et al., 2023). However, the application of LLMs in clinical diagnosis represents a more nuanced and realistic challenge, requiring not only the understanding of medical knowledge but also the ability to make complex clinical decisions based on real-life patient-specific scenarios (McDuff et al., 2023; Tu et al., 2024).
@@ 2 Related works
The literature predominantly centers on evaluations with the Focus of Medical Specialties on either a single disease such as Alzheimer’s Disease (Kwon et al., 2023), Appendicitis, Cholecystitis, Diverticulitis, and Pancreatitis (Hager et al., 2024), or a single specialty such as Ophthalmology (Lyons et al., 2023; Hu et al., 2023; Knebel et al., 2023; Rojas-Carabali et al., 2023; Madadi et al., 2023; Sorin et al., 2023), Neurology (Horiuchi et al., 2024; Galetta and Meltzer, 2023; Koga et al., 2023), and Dermatology (Stoneham et al., 2023; Rundle et al., 2024; Ravipati et al., 2023). A few works conduct their evaluation containing cases from multiple specialties, but the Dataset Size (# Test Cases) of these works is often less than 100, hardly reaching 500, resulting in limited coverage of disease types. Considering these two key features of diagnostic target scope, the current studies often narrow the disease spectrum and fail to mirror the wide breadth needed for general clinical practice.
@@
Berg et al. (2024); Brin et al. (2023); Chee et al. (2023); Delsoz et al. (2023a); Delsoz et al. (2023b); Fraser et al. (2023); Horiuchi et al. (2023a); Horiuchi et al. (2023b); Horiuchi et al. (2024); Hu et al. (2023); Koga et al. (2023); Krusche et al. (2024); Madadi et al. (2023); Mitsuyama et al. (2023); Nakaura et al. (2024); Pillai and Pillai (2023); Ravipati et al. (2023); Rojas-Carabali et al. (2023); Rundle et al. (2024); Sorin et al. (2023); Stoneham et al. (2023); Suthar et al. (2023); Tenner et al. (2024); Ueda et al. (2023a); Xv et al. (2023) Specialty-Specific $\leq 500$ Free Text Diagnoses
@@ 3.1 Realistic clinical decision tasks with expert-defined output spaces
We introduce four clinical decision tasks: deciding diagnoses, identifying procedures, ordering lab tests and prescribing medications. To reconstruct clinicians’ decision processes, different sets of information are provided as input for each task, and the target output space is also task-specific.
@@ 3.3 Balanced and diverse evaluation instances sampling
We sample a representative evaluation set from the extracted admission instances for each task, and the remaining instances can be used for training. There are no overlapped admissions and a small number of overlapped patients between split sets. We filter out the admission instances where there is no records for the target element (i.e. diagnoses, procedures, lab test orders, or prescriptions).
@@ 3.5 Matching from free text generation to medical code candidate
We design a flexible mechanism to convert LLM output to the modality required by tasks. In clinical operations, clinicians have to select from a pool of candidate codes. To avoid severe performance punishment for lack of understanding of the professional coding system, CliBench accepts natural language decisions along with codes. If the response is in codes, we use the parsed and normalized codes as predictions. If the response is in natural language, we find the most semantically similar code by calculating cosine similarity between sentence embeddings of definitions of all codes and the response produced by a BERT model trained on 1B sentence pairs (Reimers and Gurevych, 2019).
@@ 3.6 Multi-granular evaluation metrics
We compare the predicted decision set with the factual decisions and report micro precision, recall, and F1 scores across all evaluation admission instances. To reflect the LLMs’ performance in different granularities, we map each code in the predicted and ground-truth decision list to its ancestors and report the scores from coarse-grained high-level choices to fine-grained capability of distinguishing similar candidates. For diagnosis decision, we report scores for chapter-level (e.g. “endocrine, nutritional and metabolic diseases”), category group-level (e.g. “diabetes mellitus”), category-level (e.g. “diabetes mellitus due to underlying condition”), sub-category level (e.g. “Diabetes mellitus due to underlying condition with hyperosmolarity”) and full code matching (e.g. “Diabetes mellitus due to underlying condition with hyperosmolarity with coma”) following the hierarchy of ICD-10-CM. For procedure decisions, level 1 (e.g. “nuclear medicine”), level 2 (e.g. “heart-related nuclear medicine”), level 3 (e.g. “Planar Nuclear Medicine Imaging”), and full code (e.g. “Planar Nuclear Medicine Imaging of Right and Left Heart using Technetium 99m”) matching are reported. The level 1 to 3 of a full code is defined as the first 1/2/3 character(s) of the ICD-10-PCS code. For lab tests, we use levels 1 to 3 and the leaf code following LOINC hierarchy. For prescriptions, we report levels 1 to 4 of ATC classification system defined as anatomical main group, therapeutic subgroup, pharmacological subgroup and chemical subgroup. The detailed level-wise statistics can be found in Appendix B.2.
@@ 4 Comparing methods and models
Pre-trained open models includes Mistral v0.3 7B, LLaMA2 7B and LLaMA3 7B (AI@Meta, 2024). Instruction-tuned open models include three versions of Mistral Instruct 7B (Jiang et al., 2023), mixture-of-experts model Mixtral Instruct v0.1 8x7B, LLaMA2 Instruct 7B (Touvron et al., 2023), as well as 7B and 70B versions of LLaMA3 Instruct. Additionally, we include Flan-T5 XL (Chung et al., 2022) to represent models with encoder-decoder architecture. Medical specialized open models includes BioMistral DARE pretrains Mistral Instruct v0.1 on PubMed Central (Labrak et al., 2024), merging the BioMistral-7B model and the Mistral Instruct v0.1 7B model for better generalizability (Shoemake, 1985). Meditron (Chen et al., 2023) continually pretrain LLaMA2-7B on medical corpus including PubMed articles. Asclepius (Kweon et al., 2023) fine-tunes LLaMA2-7B with QA synthesized from PMC-Patients case reports (Zhao et al., 2023). OpenBioLLM (Ankit Pal, 2024) fine-tunes LLaMA3 8B model on custom medical instruct and ranking datasets. Close models includes GPT-3.5 turbo, GPT-4 turbo and GPT-4o, which are leading models without weight access. We use Azure OpenAI service and opt out of human review of the data to prevent third parties’ data access, following the MIMIC data publisher’s suggestion and complying with the data use agreement. The model versions are the ones offered on May 25, 2024. For GPT-4o, we use the global deployment variant. Fine-tuned method includes LLaMA3 Instruct 8B models after supervised fine-tuning (SFT) using LoRA adapters (Hu et al., 2021) on the training set of CliBench consisting of ground-truth sequence-to-sequence data instances.
5 Performance of LLMs on clinical decisions
Table 3: Performance on diagnosis decisions. We use zero-shot prompting for rows 1-18; we fine-tune the model on diagnosis decision training data for row 19. The “[$n$]” notation following the model name indicated that the model is fine-tuned from the model in row $n$. indicates the model is trained on biomedical or clinical resources. Bold marks the best model in each model group and green background indicates the best overall model. We report the F1 score for each level and precision, recall, and F1 score for average performance across all levels. The full result is at Appendix C.2.
@@ 5.1 Discharge diagnoses
3) GPT-4o is better than open models. The best close model (row 18) is better than the best open model (row 15), especially for fine-grained levels with a 7.37-point gap for full code. LLaMA3 Instruct 70B outperforms GPT-4 turbo.
4) Domain-specialized models do not work. Models pretrained or fine-tuned on medical or clinical corpus or tasks do not outperform (rows 6 vs 3) or slightly improve (rows 10/11 vs 8, rows 14 vs 12) its base model, but they are falling behind the instruction-tuned model in general domain. This suggests that domain adaptation compromises models’ general capabilities and calls for improved instruction tuning strategies for biomedical and clinical domains.
Table 4: Performance on procedure decisions, lab test orders and prescriptions using different abstraction levels (F1 score, %). We show full results in Appendix C.3, C.4 and C.5.
@@
13 LLaMA3 Instruct 8B 30.56 14.09 5.73 0.47 99.86 77.15 48.98 18.78 75.43 72.60 48.52 39.63
@@ 5.2 Procedure, lab orders and prescriptions
2) No emergent capabilities while scaling up for procedures and lab test orders. For diagnoses and prescriptions, we observe better or larger general domain models produce better performance (row 18 vs 16, 15 vs 13, 7 vs 3). Increasing the LLaMA3 Instruct size from 8B to 70B leads to 11.82 and 5.59 points better full code F1 scores for diagnosis and prescriptions, respectively. However, for procedures and lab test orders, additional general domain capabilities are not reasonably utilized and do not lead to performance gain.
@@ 6 Analysis
We analyze the diagnosis performance with different clinical data attributes by comparing the best-performing close model, GPT-4o, with the leading open model, LLaMA3 Instruct 70B. We provide additional analysis on the performance of various diversity and length of output in Appendix C.1.
@@ A.1 Task formulations and benchmark design
Data leak during pre-training might compromise the fairness of the benchmark. We acknowledge that it is possible that the clinical notes used for this benchmark might be used for training LLMs. We would like to emphasize that even if the model is trained with clinical notes of the same distribution, simply remembering the patterns without in-depth reasoning does not work well. As shown by the SFT results in Table 9, even when training the model with the training set of CliBench, the performance improvement is marginal. The challenging nature of CliBench makes it more robust against data leaks.
A.2 Implications and impact.

### v2

# CliBench: A Multifaceted and Multigranular Evaluation of Large Language Models for Clinical Decision Making

@@ Abstract
The integration of Artificial Intelligence (AI), especially Large Language Models (LLMs), into the clinical diagnosis process offers significant potential to improve the efficiency and accessibility of medical care. While LLMs have shown some promise in the medical domain, their application in clinical diagnosis remains underexplored, especially in real-world clinical practice, where highly sophisticated, patient-specific decisions need to be made. Current evaluations of LLMs in this field are often narrow in scope, focusing on specific diseases or specialties and employing simplified diagnostic tasks. To bridge this gap, we introduce CliBench, a novel benchmark developed from the MIMIC IV dataset, offering a comprehensive and realistic assessment of LLMs’ capabilities in clinical diagnosis. This benchmark not only covers diagnoses from a diverse range of medical cases across various specialties but also incorporates tasks of clinical significance: treatment procedure identification, lab test ordering and medication prescriptions. Supported by structured output ontologies, CliBench enables a precise and multi-granular evaluation, offering an in-depth understanding of LLM’s capability on diverse clinical tasks of desired granularity. We conduct a zero-shot evaluation of leading LLMs to assess their proficiency in clinical decision-making. Our preliminary results shed light on the potential and limitations of current LLMs in clinical settings, providing valuable insights for future advancements in LLM-powered healthcare.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2024; Anil et al., 2023) have opened up new possibilities. They have demonstrated remarkable capabilities in text understanding, reasoning, and generating proper responses in various domains (Ma et al., 2024b), including the medical field (Singhal et al., 2023a; b; Ma et al., 2023; Xu et al., 2023). Specifically, LLMs have been thoroughly evaluated and shown to excel in medical licensing examinations (Nori et al., 2023b; Gilson et al., 2023) and medical knowledge QA benchmarks (Singhal et al., 2023b; Nori et al., 2023a; Toma et al., 2023). However, the application of LLMs in clinical diagnosis represents a more nuanced and realistic challenge, requiring not only the understanding of medical knowledge but also the ability to make complex clinical decisions based on real-life patient-specific scenarios (McDuff et al., 2023; Tu et al., 2024).
@@ 2 Related works
The literature predominantly centers on evaluations with the Focus of Medical Specialties on either a single disease such as Alzheimer’s Disease (Kwon et al., 2023), Appendicitis, Cholecystitis, Diverticulitis, and Pancreatitis (Hager et al., 2024), or a single specialty such as Ophthalmology (Lyons et al., 2023; Hu et al., 2023; Knebel et al., 2023; Rojas-Carabali et al., 2023; Madadi et al., 2023; Sorin et al., 2023), Neurology (Horiuchi et al., 2024; Galetta & Meltzer, 2023; Koga et al., 2023), and Dermatology (Stoneham et al., 2023; Rundle et al., 2024; Ravipati et al., 2023). A few works conduct their evaluation containing cases from multiple specialties, but the Dataset Size (# Test Cases) of these works is often less than 100, hardly reaching 500, resulting in limited coverage of disease types. Considering these two key features of diagnostic target scope, the current studies often narrow the disease spectrum and fail to mirror the wide breadth needed for general clinical practice.
@@
Berg et al. (2024); Brin et al. (2023); Chee et al. (2023); Delsoz et al. (2023a); Delsoz et al. (2023b); Fraser et al. (2023); Horiuchi et al. (2023a); Horiuchi et al. (2023b); Horiuchi et al. (2024); Hu et al. (2023); Koga et al. (2023); Krusche et al. (2024); Madadi et al. (2023); Mitsuyama et al. (2023); Nakaura et al. (2024); Pillai & Pillai (2023); Ravipati et al. (2023); Rojas-Carabali et al. (2023); Rundle et al. (2024); Sorin et al. (2023); Stoneham et al. (2023); Suthar et al. (2023); Tenner et al. (2024); Ueda et al. (2023a); Xv et al. (2023) Specialty-Specific $\leq 500$ Free Text Diagnoses
@@ 3.1 Realistic clinical decision tasks with expert-defined output spaces
We introduce four clinical decision tasks: deciding diagnoses, identifying procedures, ordering lab tests and prescribing medications. To reconstruct clinicians’ decision processes, different sets of information are provided as input for each task, and the target output space is also task-specific. We show a table laying out the comparisons and examples in Appendix B.1.
@@ 3.3 Balanced and diverse evaluation instances sampling
We sample a representative evaluation set from the extracted admission instances for each task, and the remaining instances can be used for training. There are no overlapped admissions or patients between split sets. We filter out the admission instances where there is no records for the target element (i.e. diagnoses, procedures, lab test orders, or prescriptions).
@@ 3.5 Matching from free text generation to medical code candidate
We design a flexible mechanism to convert LLM output to the modality required by tasks. In clinical operations, clinicians have to select from a pool of candidate codes. To avoid severe performance punishment for lack of understanding of the professional coding system, CliBench accepts natural language decisions along with codes. If the response is in codes, we use the parsed and normalized codes as predictions. If the response is in natural language, we find the most semantically similar code by calculating cosine similarity between sentence embeddings of definitions of all codes and the response produced by a BERT model trained on 1B sentence pairs (Reimers & Gurevych, 2019).
@@ 3.6 Multi-granular evaluation metrics
We compare the predicted decision set with the factual decisions and report micro precision, recall, and F1 scores across all evaluation admission instances. To reflect the LLMs’ performance in different granularities, we map each code in the predicted and ground-truth decision list to its ancestors and report the scores from coarse-grained high-level choices to fine-grained capability of distinguishing similar candidates. The task is more difficult for finer-grained granularities with a larger candidate pool. LLMs might be able to predict the correct disease chapters out of 21 chapters correctly but struggle to decide the specific disease category groups with 283 possibilities.
For diagnosis decision, we report scores for chapter-level (e.g. “endocrine, nutritional and metabolic diseases”), category group-level (e.g. “diabetes mellitus”), category-level (e.g. “diabetes mellitus due to underlying condition”), sub-category level (e.g. “Diabetes mellitus due to underlying condition with hyperosmolarity”) and full code matching (e.g. “Diabetes mellitus due to underlying condition with hyperosmolarity with coma”) following the hierarchy of ICD-10-CM. For procedure decisions, level 1 (e.g. “nuclear medicine”), level 2 (e.g. “heart-related nuclear medicine”), level 3 (e.g. “Planar Nuclear Medicine Imaging”), and full code (e.g. “Planar Nuclear Medicine Imaging of Right and Left Heart using Technetium 99m”) matching are reported. The level 1 to 3 of a full code is defined as the first 1/2/3 character(s) of the ICD-10-PCS code. For lab tests, we use levels 1 to 3 and the leaf code following LOINC hierarchy. For prescriptions, we report levels 1 to 4 of ATC classification system defined as anatomical main group, therapeutic subgroup, pharmacological subgroup and chemical subgroup. The statistics of the candidate decision space sizes of all levels shown in Appendix B.3 demonstrate the detail level and difficulty of each granularity.
@@ 4 Comparing methods and models
Pre-trained open models includes Mistral v0.3 7B, LLaMA2 7B and LLaMA3 8B (AI@Meta, 2024). Instruction-tuned open models include three versions of Mistral Instruct 7B (Jiang et al., 2023), mixture-of-experts model Mixtral Instruct v0.1 8x7B, LLaMA2 Instruct 7B (Touvron et al., 2023), 8B and 70B versions of LLaMA3 Instruct, LLaMA3.1 Instruct 8B, as well as 1B and 3B versions of LLaMA3.2 Instruct. Additionally, we include Flan-T5 XL (Chung et al., 2022) to represent models with encoder-decoder architecture. Medical specialized open models includes BioMistral DARE pretrains Mistral Instruct v0.1 on PubMed Central (Labrak et al., 2024), merging the BioMistral-7B model and the Mistral Instruct v0.1 7B model for better generalizability (Shoemake, 1985). Meditron (Chen et al., 2023) continually pretrain LLaMA2-7B on medical corpus including PubMed articles. Asclepius (Kweon et al., 2023) fine-tunes LLaMA2-7B with QA synthesized from PMC-Patients case reports (Zhao et al., 2023). OpenBioLLM (Ankit Pal, 2024) fine-tunes LLaMA3 8B model on custom medical instruct and ranking datasets. Close models includes GPT-3.5 turbo, GPT-4 turbo and GPT-4o, which are leading models without weight access. We use Azure OpenAI service and opt out of human review of the data to prevent third parties’ data access, following the MIMIC data publisher’s suggestion and complying with the data use agreement. The model versions are the ones offered on May 25, 2024. For GPT-4o, we use the global deployment variant. Fine-tuned method is the LLaMA3 Instruct 8B models after supervised fine-tuning (SFT) using LoRA adapters (Hu et al., 2021) on the training set of CliBench consisting of ground-truth sequence-to-sequence instances.
Table 3: Performance on diagnosis decisions. We use zero-shot prompting for rows 1-21; we fine-tune the model on diagnosis decision training data for row 22. The “[$n$]” notation following the model name indicated that the model is fine-tuned from the model in row $n$. indicates the model is trained on biomedical or clinical resources. Bold marks the best model in each model group and green background indicates the best overall model. We report the F1 score for each level and precision, recall, and F1 score for average performance across all levels. The full result is at Appendix C.1.
@@
14 LLaMA3.1 Instruct 8B 63.66 39.20 26.48 13.71 9.97 27.76 34.14 30.61
15 LLaMA3.2 Instruct 1B 55.19 28.98 16.82 7.95 5.34 23.92 21.88 22.86
16 LLaMA3.2 Instruct 3B 61.26 33.90 20.85 10.14 6.79 25.23 28.13 26.59
@@ 5.1 Discharge diagnoses
3) GPT-4o is better than open models. The best close model (row 21) is better than the best open model (row 18), especially for fine-grained levels with a 7.37-point gap for full code. LLaMA3 Instruct 70B outperforms GPT-4 turbo.
4) Domain-specialized models do not work. Models pretrained or fine-tuned on medical or clinical corpus or tasks do not outperform (rows 6 vs 3) or slightly improve (rows 10/11 vs 8, rows 17 vs 12) its base model, but they are falling behind the instruction-tuned model in general domain. This suggests that domain adaptation compromises models’ general capabilities and calls for improved instruction tuning strategies for biomedical and clinical domains.
Table 4: Performance on procedure decisions, lab test orders and prescriptions using different abstraction levels (F1 score, %). We show full results in Appendix C.2, C.3 and C.4.
@@
13 LLaMA3 Instruct 8B 30.56 14.09 5.73 0.39 99.86 77.15 48.98 18.78 75.43 72.60 48.52 39.63
14 LLaMA3.1 Instruct 8B 28.98 13.15 5.24 0.39 99.91 77.40 48.67 19.86 79.32 76.74 50.50 40.26
15 LLaMA3.2 Instruct 1B 25.79 11.57 4.26 0.26 99.86 76.43 44.82 15.90 73.34 69.78 43.44 33.99
16 LLaMA3.2 Instruct 3B 27.99 14.37 5.14 0.24 99.91 77.20 48.38 21.83 76.05 72.97 45.10 34.20
@@ 5.2 Procedure, lab orders and prescriptions
2) No emergent capabilities while scaling up for procedures and lab test orders. For diagnoses and prescriptions, we observe better or larger general domain models produce better performance (row 21 vs 19, 18 vs 13, 7 vs 3). Increasing the LLaMA3 Instruct size from 8B to 70B leads to 11.82 and 5.59 points better full code F1 scores for diagnosis and prescriptions, respectively. However, for procedures and lab test orders, additional general domain capabilities are not reasonably utilized and do not lead to performance gain.
@@ 6 Analysis
We analyze the diagnosis performance with different clinical data attributes by comparing the best-performing close model, GPT-4o, with the leading open model, LLaMA3 Instruct 70B. We provide additional analysis on the performance of various diversity and length of output in Appendix 6.2.
@@ 6.1 Diagnosis capabilities by patient attributes
6.2 Diagnosis capabilities by task difficulties
We demonstrate breakdown performance for admissions with various durations in Figure 3(a). Both GPT-4o and LLaMA3 70B Instruct models produce worse recall when the duration increases, as the longer the stay, the more complicated the diagnosis decisions to make. In Figure 3(b), we analyze the performance of admission subsets with different diversity of diseases. The number of unique chapters represents the diversity and scope of diagnosis involved in the ground-truth billing code. The precision increases while a broader range of unique chapters is involved. In Figure 3(c), we show the performance trend for admission instances with different numbers of ground-truth diagnoses. With the increase in number of diagnoses, the recall is getting lower while the precision is getting higher. From qualitative examples, we observe the models generate similar amounts of predictions no matter the complexity of the patient case. With a larger number of ground-truth diagnoses, it is easier to have matches, leading to improved precision.
Figure 3: The breakdown diagnosis performance for full code level by the difficulties of the patient cases, specifically: (a) how long the patient stayed in the hospital, (b) the number of unique diagnosis chapters in ground-truth billing diagnosis codes, (c) the number of diagnosis codes at discharge.
@@ A.1 Task formulations and benchmark design
Data from a single medical center may limit the generalizability of the findings. 1) To create a publicly accessible representative clinical benchmark with reasonable size, we have to make tradeoff choices. Given the limited test set size (to avoid the evaluation being too heavy), we prioritize the diversity of output spaces (such as 21 diagnosis chapters), 21 service departments and 37 care units. This ensures the evaluation covers a wide range of clinical scenarios and outcomes. We deprioritized sampling from different medical centers from additional data sources. 2) Though having more medical centers would make CliBench even more diverse, we would like to reiterate that CliBench already demonstrates a much stronger generalizability containing patient cases from multiple specialties with a much larger expert ontology as output space, and supports four tasks spanning the clinical operation lifecycles as we thoroughly compared in Table 1. 3) We acknowledge that our proposed benchmark construction method can be applied to an even more comprehensive range of data sources.
@@
Data leak during pre-training might compromise the fairness of the benchmark. 1) The license of the data source used for CliBench requires that models trained on the data should be treated as containing sensitive information and thus have to inherit the agreement of MIMIC-IV. This requirement ensures that models trained on the data have to disclose the source, making data leakage easy to recognize. This also means that all models that are open to download without signing an agreement (such as most models on Huggingface Hub) should not use MIMIC data for training.
2) The challenging nature of CliBench makes it more robust against data leaks. Unlike many existing benchmarks on factual questions with simpler reasoning processes, we show that even if the model is fine-tuned with the clinical notes of the same distribution, simply remembering the patterns without in-depth reasoning does not perform well (as shown by the SFT performance in Table 3).
3) We acknowledge in the limitation section (Appendix D) that clinical notes used for CliBench might be used for training LLMs and the risk of data leakage. This issue is inevitable for LLM benchmarks using public-accessible data sources.
Generalizability of the training data might be limited. 1) We do not claim that any of our evaluated methods or the training data has a strong generalizability toward novel medical conditions. We do not focus on training a model for this task. We claim that we provide a more realistic and inclusive benchmark of LLM’s clinical decision capabilities than existing benchmarks. As we mentioned in the “Limitationn” section in Appendix D, better generalizability with even more data or method design is possible. 2) In our SFT performance shown in Table 10, we observed that fine-tuning on the training data can only provide marginal improvement compared with zero-shot. However, this does not undermine the value of the training data. Instead, it calls for better method design and utilization strategy for the training data.
Heavy dependence on extensive manual checking. 1) We would like to clarify that CliBench does NOT manually check the model responses to produce the evaluation scores. Instead, we develop an automatic evaluation strategy to produce evaluation metrics from LM’s free-text response in a scalable way, as introduced in Section 3.5 and Section 3.6. 2) We use manual checking only when creating the data processing pipeline and constructing the input prompt of evaluation test cases (as discussed in Section 3.2).
Why not provide the decision of a doctor as the upper bound performance? 1) We use the clinical decisions recorded in the raw EHR data as the ground truth for the four clinical decision tasks. These decisions are made by different specialized clinicians in the corresponding service department or unit. We then consider these recorded decisions as ground truth, representing the “collective knowledge” of a group of clinicians.
2) As we acknowledge in the “Limitations” section in Appendix D, there could be miscoded decisions, however, the recorded diagnoses/procedures/lab test orders/prescriptions are the best resources for clinical decision labels we can get from existing available data.
3) Since CliBench covers patient cases from many different units, annotations from a small group of doctors are not sufficient to have reasonable coverage. Obtaining doctors’ performance from various units has a non-trivial cost.
A.2 Implications and impact